# Real-world models for multiple term structures: a unifying HJM semimartingale framework

### v2

@@ 3.2. Characterization of LMDs
[…]
where $\mathcal{X}_{t,T}^{i}:=\{x\in E:c^{i}_{t}(x)>2e^{\bar{\gamma}^{i}(t,T,x)}-1\}$, and the following two conditions hold a.s.:
@@
Moreover, a strictly positive local martingale $Z=(Z_{t})_{t\geq 0}$ belongs to $\mathcal{D}$ if and only if
(3.6) $Z=\mathcal{E}\left(\lambda\cdot W+\psi\ast\tilde{\mu}+N\right),$
@@
(i) Let us consider the function $h:\mathbb{R}\to[0,1]$ given by $h(x):=x\mathbf{1}_{\{|x|\leq 1\}}$. By [35, Theorem II.2.34], the canonical representation of the semimartingale $Y$ corresponding to the truncation function $h$ is
@@
so that the finite variation process $(v-h\circ(u+v))\ast\mu$ is a special semimartingale and, therefore, of locally integrable variation (see, e.g., [35, Proposition I.4.23]). In turn, this implies that the process $|v-h\circ(u+v)|\ast\nu$ is locally integrable and, hence, equation (3.9) can be rewritten more simply as
@@
(ii) Let $Z\in\mathcal{D}$. Since $Z$ is a strictly positive local martingale with $Z_{0}=1$, its stochastic logarithm $L:=Z^{-1}_{-}\cdot Z$ is a local martingale with $L_{0}=0$ and $\Delta L>-1$. By [35, Lemma III.4.24], it holds that
@@
for some $\lambda\in L^{2}_{\rm loc}(W)$, $\psi\in G_{\rm loc}(\mu)$ and a local martingale $N$ with $N_{0}=0$ satisfying $\langle N,W^{j}\rangle=0$, for all $j=1,\ldots,d$, and $M_{\mu}[\Delta N|\widetilde{\mathcal{P}}]=0$. Moreover, by part b) of [35, Theorem III.4.20], it holds that $\psi=M_{\mu}[\Delta L|\widetilde{\mathcal{P}}]$. Since $\Delta L>-1$, this directly implies that the function $\psi$ takes values in $(-1,+\infty)$. By Lemma 3.3 and an application of Yor’s formula (see, e.g., [35, formula II.8.19]) we obtain that
$Z(X^{0})^{-1}S^{i}B^{i}(\cdot,T)=S^{i}_{0}B^{i}(0,T)\mathcal{E}(L+Y+[L,Y]).$
This shows that $Z\in\mathcal{D}$ if and only if $Y+[L,Y]$ is a local martingale. From (3.8) and (3.11) we obtain
$[L,Y]=\int_{0}^{\cdot}\lambda^{\top}_{s}\sigma_{s}{\rm d}s+\sum_{s>0}\Delta L_{s}\Delta Y_{s}=\int_{0}^{\cdot}\lambda^{\top}_{s}\sigma_{s}{\rm d}s+\bigl{(}(\psi+\Delta N)(u+v)\bigr{)}\ast\mu.$
@@
where we made use of the $\widetilde{\mathcal{P}}$-measurability of the functions $u$, $v$ and $\psi$ together with the fact that $M_{\mu}[\Delta N|\widetilde{\mathcal{P}}]=0$. By compensating $C(h)$, we obtain that the process in (3.12) is a local martingale if and only if the process $D:=B(h)+\int_{0}^{\cdot}\lambda^{\top}_{s}\sigma_{s}{\rm d}s+C(h)^{p}$ is a local martingale. By (3.10), the process $D$ can be equivalently rewritten as
(3.14) $D=A+\int_{0}^{\cdot}\lambda^{\top}_{s}\sigma_{s}{\rm d}s+\bigl{(}(u+v)\psi+u\bigr{)}\ast\nu.$
Since $D$ is a predictable process of finite variation, it can be a local martingale if and only if it is equal to zero, up to an evanescent set (see [35, Corollary I.3.16]). Recalling the notation introduced at the beginning of the proof, this means that, outside of a subset of $\Omega\times\mathbb{R}_{+}$ of $(\mathbb{P}\otimes{\rm d}t)$-measure zero,
@@
(iii) To complete the first part of the proof, it remains to show that (3.5) holds. Since the compensator $C(h)^{p}$ introduced above is a predictable process of finite variation, its variation
@@
$\bigl{(}(u+v)(1+\psi)\mathbf{1}_{\{u+v>1\}}\bigr{)}\ast\nu=\int_{0}^{\cdot}\int_{\mathcal{X}^{i}_{s,T}}\bigl{(}(1+c^{i}_{s}(x))e^{-\bar{\gamma}^{i}(s,T,x)}-1\bigr{)}\bigl{(}1+\psi_{s}(x)\bigr{)}F_{s}({\rm d}x){\rm d}s$
is locally integrable, where the set $\mathcal{X}^{i}_{s,T}\subset E$ has been defined as in the statement of theorem. This proves the validity of the integrability condition (3.5) for a.e. $t\in[0,T]$ whenever $\mathcal{D}\neq\emptyset$.
(iv) Conversely, suppose there exist $\lambda\in L^{2}_{\rm loc}(W)$ and $\psi\in G_{\rm loc}(\mu)$ with $\psi>-1$ such that (3.5) and conditions (i)-(ii) in the statement of the theorem hold. We shall prove that $\mathcal{D}\neq\emptyset$. To this effect, let us define a strictly positive local martingale $Z$ as in (3.6), i.e.,
$Z:=\mathcal{E}(L),\qquad\text{ with }L:=\lambda\cdot W+\psi\ast\tilde{\mu}+N,$
@@
The process $\mathbf{1}_{\{|\Delta Y|\leq 1\}}\cdot[Y]$ is locally bounded, being an increasing process with bounded jumps, while the process $(\sum_{s>0}\Delta L_{s}^{2})^{1/2}$ is locally integrable, due to the fact that $L$ is a local martingale (see, e.g., [35, Corollary I.4.55]). This implies that the increasing process $|h\circ(u+v)(\psi+\Delta N)|\ast\mu$ is locally integrable. We then proceed by showing that the increasing process $|\mathbf{1}_{\{u+v>1\}}(u+v)(1+\psi+\Delta N)|\ast\mu$ is locally integrable too. Notice that $u+v\geq-1$ and also $01+\psi+\Delta N>$ 0, since $\Delta L>-1$. Therefore, by Lemma B.1 and using the fact that $M_{\mu}[\Delta N|\widetilde{\mathcal{P}}]=0$, the process $(\mathbf{1}_{\{u+v>1\}}(u+v)(1+\psi+\Delta N))\ast\mu$ is locally integrable if and only if $(\mathbf{1}_{\{u+v>1\}}(u+v)(1+\psi))\ast\nu$ is locally integrable. To prove that the latter property holds, observe first that (3.5) corresponds to the following condition, for a.e. $t\in[0,T]$:
@@
where we have denoted by $\dot{A}$ and $\dot{B}(h)$ the densities of the absolutely continuous processes $A$ and $B(h)$ with respect to the Lebesgue measure, respectively. In the second equality of (3.17), we made use of the fact that the increasing process $|v-h\circ(u+v)|\ast\nu$ is locally integrable and, therefore, the integral $\int_{E}(v_{t}(x)-h(u_{t}(x)+v_{t}(x)))F_{t}({\rm d}x)$ is always finite outside a set of $(\mathbb{P}\otimes{\rm d}t)$-measure zero. Since $\lambda$ and $\sigma$ belong to $L^{2}_{\rm loc}(W)$, the Cauchy-Schwarz inequality implies that $\int_{0}^{T}|\lambda^{\top}_{t}\sigma_{t}|{\rm d}t<+\infty$ a.s. Moreover, the same estimates as in (3.16) together with an application of Lemma B.1 allow to show that $(h\circ(u+v)\psi)\ast\nu$ is well-defined as a predictable process of finite variation. In view of (3.17), these facts enable us to deduce that $(\mathbf{1}_{\{u+v>1\}}(u+v)(1+\psi))\ast\nu$ is well-defined as an increasing process. Since every predictable increasing process is locally integrable (see, e.g., [35, Lemma I.3.10]) and
@@
we have thus proved that the finite variation process $C(h)$ is of locally integrable variation. Therefore, its compensator $C(h)^{p}$ exists and is given by (3.13). Since conditions (i)-(ii) together imply that the process $D$ in (3.14) vanishes, the process (3.12) results to be a local martingale. As explained above, the latter property is equivalent to the local martingale property of $Y+[L,Y]$. Since $Z$ is strictly positive, this suffices to conclude that $Z\in\mathcal{D}$, thus completing the proof. ∎
@@
Theorem 3.4 implies that, whenever $\mathcal{D}\neq\emptyset$, it always holds that $\widehat{Z}:=\mathcal{E}(\lambda\cdot W+\psi\ast\tilde{\mu})\in\mathcal{D}$, where $\lambda$ and $\psi$ satisfy (3.5) and conditions (i)-(ii) of the theorem. The LMD $\widehat{Z}$ corresponds to taking $N\equiv 0$ in (3.6) and can be regarded as the minimal LMD, in the spirit of the minimal martingale measure introduced in [27] (see also [15, Corollary 2.14] in a finite-dimensional semimartingale setup).
@@
provided that $\int_{E}\sup_{T\geq t}|\gamma^{0}(t,T,x)(1-(1+\psi_{t}(x))\exp(-\bar{\gamma}^{0}(t,T,x)))|F_{t}({\rm d}x)<+\infty$, so that we are allowed to differentiate under the integral sign. This condition represents the real-world HJM drift restriction for the riskless term structure and has been already derived in [13] and [47, Section 3], albeit under more stringent technical requirements. For $i\in I$, differentiating condition (ii) leads to the real-world HJM drift restriction for the risky term structures:
@@ 3.3. Monotonicity of term structures
The examples discussed in Section 1.1 show that, in many cases, there exists a natural ordering among the risky term structures according to their level of risk. It is, therefore, of interest to provide clear conditions ensuring that, for each $T>0$, the family $\{S^{i}B^{i}(\cdot,T);i\in I\}$ is ordered with respect to $i$. Here, we do not aim at a general characterization of ordered term structures, but rather at simple criteria that can be easily verified in concrete specification. To this end, we introduce the following condition.
@@
Let $Z\in\mathcal{D}$ and $i\in I_{0}$. We say that the term structure indexed by $i$ is fairly priced by $Z$ if the process $Z(X^{0})^{-1}S^{i}B^{i}(\cdot,T)$ is a true martingale, for every $T>0$.
@@
In view of Definition 2.4, the process $Z(X^{0})^{-1}S^{i}B^{i}(\cdot,T)$ is a local martingale, for every $Z\in\mathcal{D}$ and $(i,T)\in I_{0}\times\mathbb{R}_{+}$. Definition 3.8 strengthens this property into a true martingale property. This corresponds to a generalization of the concept of fair pricing in the context of the benchmark approach, where a price process is said to be fair if it is a true martingale when denominated in units of the growth-optimal portfolio (see [46, Chapter 10] and [13, Section 2.2] for term structure models).
@@
Suppose that Assumption 3.1 holds and $\mathcal{D}\neq\emptyset$. For $Z\in\mathcal{D}$ and $i,j\in I$ with $i<j$, assume that the term structures indexed by $i$ and $j$ are fairly priced by $Z$, in the sense of Definition 3.8. Assume furthermore that Condition 3.7 holds and $S^{i}_{0}\leq S^{j}_{0}$. Then, it holds a.s. that
@@
The assumptions imply that the process $Z(X^{0})^{-1}S^{k}B^{k}(\cdot,T)$ is a true martingale, for $k\in\{i,j\}$ and for all $T>0$. Therefore, making use of the terminal condition $B^{k}(T,T)=1$, it follows that
[…]
@@ 3.4. The risk-neutral setup
We have so far developed a general framework which does not rely on the existence of a risk-neutral probability measure, being based on a weaker concept of market viability (see Section 2.2). Given the widespread use of risk-neutral valuation in finance, it is nevertheless useful to specialize Theorem 3.4 to the risk-neutral setup. Adopting the notation of Section 2.2, we say that a probability measure $\mathbb{Q}\sim\mathbb{P}$ on $(\Omega,\mathcal{F})$ is a risk-neutral probability if $X^{(i,T)}$ is a $\mathbb{Q}$-local martingale, for every $(i,T)\in I_{0}\times\mathbb{R}_{+}$. We recall from Girsanov’s theorem (see [35, Theorems III.3.11 and III.3.17]) that, if $\mathbb{Q}\sim\mathbb{P}$, there exist $\lambda\in L^{2}_{\rm loc}(W)$ and $\psi\in G_{\rm loc}(\mu)$ with $\psi>-1$ such that $W^{\mathbb{Q}}:=W+\int_{0}^{\cdot}\lambda_{s}{\rm d}s$ is a $d$-dimensional Brownian motion under $\mathbb{Q}$ and the compensator of $\mu$ under $\mathbb{Q}$ is given by
@@
Let $\mathbb{Q}$ be a probability measure on $(\Omega,\mathcal{F})$ with $\mathbb{Q}\sim\mathbb{P}$ and density process $Z=(Z_{t})_{t\geq 0}$, i.e., $Z_{t}={\rm d}\mathbb{Q}|_{\mathcal{F}_{t}}/{\rm d}\mathbb{P}|_{\mathcal{F}_{t}}$, for all $t\geq 0$. [35, Lemma III.5.17] implies that $Z$ admits a representation of the form (3.6). By definition, $\mathbb{Q}$ is a risk-neutral probability if and only if $Z\in\mathcal{D}$. In view of Theorem 3.4, we only have to prove that (3.19)–(3.22) hold under $\mathbb{Q}$. For every $i\in I$, the process $Z^{i}$ can be represented as in (3.19) if and only if it is a special semimartingale under $\mathbb{Q}$ and this holds if and only if $c^{i}\mathbf{1}_{\{c^{i}>1\}}\ast\nu^{\mathbb{Q}}$ is of finite variation. Condition (3.5) for $t=T$ implies that $\int_{\{x\in E:c^{i}_{t}(x)>1\}}c^{i}_{t}(x)F^{\mathbb{Q}}_{t}({\rm d}x)<+\infty$ a.s. for a.e. $t\in\mathbb{R}_{+}$. Making use of condition (i) in Theorem 3.4 and the fact that $Z^{i}$ is a special semimartingale under $\mathbb{P}$, as a consequence of (3.1), we can compute
@@
    When applied to foreign exchange models (see Section 1.1.1), Corollary 3.12 extends the result of [42, Proposition 2.1.15] by requiring substantially weaker integrability properties.

### v3

@@ 3.2. Characterization of LMDs
(3.5) $\int_{\mathcal{C}_{t,T}^{i}}\bigl((1+c^{i}_{t}(x))e^{-\bar{\gamma}^{i}(t,T,x)}-1\bigr)\bigl(1+\psi_{t}(x)\bigr)F_{t}({\rm d}x)<+\infty\quad\text{a.s.},$
where $\mathcal{C}_{t,T}^{i}:=\{x\in E:c^{i}_{t}(x)>2e^{\bar{\gamma}^{i}(t,T,x)}-1\}$, and the following two conditions hold a.s.:
@@
Moreover, a strictly positive local martingale $D=(D_{t})_{t\geq 0}$ belongs to $\mathcal{D}$ if and only if
(3.6) $D=\mathcal{E}\left(\lambda\cdot W+\psi\ast\tilde{\mu}+N\right),$
@@
(1) Let us consider the function $h:\mathbb{R}\to[0,1]$ given by $h(x):=x\mathbf{1}_{\{|x|\leq 1\}}$. By [37, Theorem II.2.34], the canonical representation of the semimartingale $Y$ corresponding to the truncation function $h$ is
@@
so that the finite variation process $(v-h\circ(u+v))\ast\mu$ is a special semimartingale and, therefore, of locally integrable variation (see, e.g., [37, Proposition I.4.23]). In turn, this implies that the process $|v-h\circ(u+v)|\ast\nu$ is locally integrable and, hence, equation (3.9) can be rewritten more simply as
@@
(2) Let $D\in\mathcal{D}$. Since $D$ is a strictly positive local martingale with $D_{0}=1$, its stochastic logarithm $L:=D^{-1}_{-}\cdot D$ is a local martingale with $L_{0}=0$ and $\Delta L>-1$. By [37, Lemma III.4.24], it holds that
@@
for some $\lambda\in L^{2}_{\rm loc}(W)$, $\psi\in G_{\rm loc}(\mu)$ and a local martingale $N$ with $N_{0}=0$ satisfying $\langle N,W^{j}\rangle=0$, for all $j=1,\ldots,d$, and $M_{\mu}[\Delta N|\widetilde{\mathcal{P}}]=0$. Moreover, by part b) of [37, Theorem III.4.20], it holds that $\psi=M_{\mu}[\Delta L|\widetilde{\mathcal{P}}$]. Since $\Delta L>-1$, this directly implies that the function $\psi$ takes values in $(-1,+\infty)$. By Lemma 3.3 and an application of Yor’s formula (see, e.g., [37, formula II.8.19]) we obtain that
$D(X^{0})^{-1}S^{i}B^{i}(\cdot,T)=S^{i}_{0}B^{i}(0,T)\mathcal{E}(L+Y+[L,Y]).$
This shows that $D\in\mathcal{D}$ if and only if $Y+[L,Y]$ is a local martingale. From (3.8) and (3.11) we obtain
$L,Y]=\int_{0}^{\cdot}\lambda^{\top}_{s}\sigma_{s}{\rm d}s+\sum_{s>0}\Delta L_{s}\Delta Y_{s}=\int_{0}^{\cdot}\lambda^{\top}_{s}\sigma_{s}{\rm d}s+\bigl((\psi+\Delta N)(u+v)\bigr)\ast\mu.$
@@
where we made use of the $\widetilde{\mathcal{P}}-$measurability of the functions $u$, $v$ and $\psi$ together with the fact that $M_{\mu}[\Delta N|\widetilde{\mathcal{P}}]=0$. By compensating $C(h)$, we obtain that the process in (3.12) is a local martingale if and only if the process $H:=B(h)+\int_{0}^{\cdot}\lambda^{\top}_{s}\sigma_{s}{\rm d}s+C(h)^{p}$ is a local martingale. By (3.10), the process $H$ can be equivalently rewritten as
(3.14) $H=A+\int_{0}^{\cdot}\lambda^{\top}_{s}\sigma_{s}{\rm d}s+\bigl((u+v)\psi+u\bigr)\ast\nu.$
Since $H$ is a predictable process of finite variation, it can be a local martingale if and only if it is equal to zero, up to an evanescent set (see [37, Corollary I.3.16]). Recalling the notation introduced at the beginning of the proof, this means that, outside of a subset of $\Omega\times\mathbb{R}_{+}$ of $(\mathbb{P}\otimes{\rm d}t)$-measure zero,
@@
(3) To complete the first part of the proof, it remains to show that (3.5) holds. Since the compensator $C(h)^{p}$ introduced above is a predictable process of finite variation, its variation
@@
$\bigl((u+v)(1+\psi)\mathbf{1}_{\{u+v>1\}}\bigr)\ast\nu=\int_{0}^{\cdot}\int_{\mathcal{C}^{i}_{s,T}}\bigl((1+c^{i}_{s}(x))e^{-\bar{\gamma}^{i}(s,T,x)}-1\bigr)\bigl(1+\psi_{s}(x)\bigr)F_{s}({\rm d}x){\rm d}s$
is locally integrable, where the set $\mathcal{C}^{i}_{s,T}\subset E$ has been defined as in the statement of theorem. This proves the validity of the integrability condition (3.5) for a.e. $t\in[0,T]$ whenever $\mathcal{D}\neq\emptyset$.
(4) Conversely, suppose there exist $\lambda\in L^{2}_{\rm loc}(W)$ and $\psi\in G_{\rm loc}(\mu)$ with $\psi>-1$ such that (3.5) and conditions (i)-(ii) in the statement of the theorem hold. We shall prove that $\mathcal{D}\neq\emptyset$. To this effect, let us define a strictly positive local martingale $D$ as in (3.6), i.e.,
$D:=\mathcal{E}(L),\qquad\text{ with }L:=\lambda\cdot W+\psi\ast\tilde{\mu}+N,$
@@
The process $\mathbf{1}_{\{|\Delta Y|\leq 1\}}\cdot[Y]$ is locally bounded, being an increasing process with bounded jumps, while the process $(\sum_{s>0}\Delta L_{s}^{2})^{1/2}$ is locally integrable, due to the fact that $L$ is a local martingale (see, e.g., [37, Corollary I.4.55]). This implies that the increasing process $|h\circ(u+v)(\psi+\Delta N)|\ast\mu$ is locally integrable. We then proceed by showing that the increasing process $|\mathbf{1}_{\{u+v>1\}}(u+v)(1+\psi+\Delta N)|\ast\mu$ is locally integrable too. Notice that $u+v\geq-1$ and also $1+\psi+\Delta N>0$, since $\Delta L>-1$. Therefore, by Lemma B.1 and using the fact that $M_{\mu}[\Delta N|\widetilde{\mathcal{P}}]=0$, the process $(\mathbf{1}_{\{u+v>1\}}(u+v)(1+\psi+\Delta N))\ast\mu$ is locally integrable if and only if $(\mathbf{1}_{\{u+v>1\}}(u+v)(1+\psi))\ast\nu$ is locally integrable. To prove that the latter property holds, observe first that (3.5) corresponds to the following condition, for a.e. $t\in[0,T]$:
@@
where we have denoted by $\dot{A}$ and $\dot{B}(h)$ the densities of the absolutely continuous processes $A$ and $B(h)$ with respect to the Lebesgue measure, respectively. In the second equality of (3.17), we made use of the fact that the increasing process $|v-h\circ(u+v)|\ast\nu$ is locally integrable and, therefore, the integral $\int_{E}(v_{t}(x)-h(u_{t}(x)+v_{t}(x)))F_{t}({\rm d}x)$ is always finite outside a set of $(\mathbb{P}\otimes{\rm d}t)$-measure zero. Since $\lambda$ and $\sigma$ belong to $L^{2}_{\rm loc}(W)$, the Cauchy-Schwarz inequality implies that $\int_{0}^{T}|\lambda^{\top}_{t}\sigma_{t}|{\rm d}t<+\infty$ a.s. Moreover, the same estimates as in (3.16) together with an application of Lemma B.1 allow to show that $(h\circ(u+v)\psi)\ast\nu$ is well-defined as a predictable process of finite variation. In view of (3.17), these facts enable us to deduce that $(\mathbf{1}_{\{u+v>1\}}(u+v)(1+\psi))\ast\nu$ is well-defined as an increasing process. Since every predictable increasing process is locally integrable (see, e.g., [37, Lemma I.3.10]) and
@@
we have thus proved that the finite variation process $C(h)$ is of locally integrable variation. Therefore, its compensator $C(h)^{p}$ exists and is given by (3.13). Since conditions (i)-(ii) together imply that the process $H$ in (3.14) vanishes, the process (3.12) results to be a local martingale. As explained above, the latter property is equivalent to the local martingale property of $Y+[L,Y]$. Since $D$ is strictly positive, this suffices to conclude that $D\in\mathcal{D}$, thus completing the proof. ∎
@@
Theorem 3.4 implies that, whenever $\mathcal{D}\neq\emptyset$, it always holds that $\widehat{D}:=\mathcal{E}(\lambda\cdot W+\psi\ast\tilde{\mu})\in\mathcal{D}$, where $\lambda$ and $\psi$ satisfy (3.5) and conditions (i)-(ii) of the theorem. The LMD $\widehat{D}$ corresponds to taking $N\equiv 0$ in (3.6) and can be regarded as the minimal LMD, in the spirit of the minimal martingale measure introduced in [27] (see also [15, Corollary 2.14] in a finite-dimensional semimartingale setup).
@@
provided that $\int_{E}\sup_{T\geq t}|\gamma^{0}(t,T,x)(1-(1+\psi_{t}(x))\exp(-\bar{\gamma}^{0}(t,T,x)))|F_{t}({\rm d}x)<+\infty$, so that we are allowed to differentiate under the integral sign. This condition represents the real-world HJM drift restriction for the riskless term structure and has already been derived in [13] and [52, Section 3], albeit under more stringent technical requirements. For $i\in I$, differentiating condition (ii) leads to the real-world HJM drift restriction for the risky term structures:
@@ 3.3. Monotonicity of term structures
The examples discussed in Section 1.1 show that, in many cases, there exists a natural ordering among the risky term structures according to their level of risk. It is, therefore, of interest to provide clear conditions ensuring that, for each $T>0$, the family $\{S^{i}B^{i}(\cdot,T);i\in I\}$ is ordered with respect to $i$. Here, we do not aim at a general characterization of ordered term structures, but rather at simple criteria that can be easily verified in concrete specifications. To this end, we introduce the following condition.
@@
Let $D\in\mathcal{D}$ and $i\in I_{0}$. We say that the term structure indexed by $i$ is fairly priced by $D$ if the process $D(X^{0})^{-1}S^{i}B^{i}(\cdot,T)$ is a true martingale, for every $T>0$.
@@
In view of Definition 2.4, the process $D(X^{0})^{-1}S^{i}B^{i}(\cdot,T)$ is a local martingale, for every $D\in\mathcal{D}$ and $(i,T)\in I_{0}\times\mathbb{R}_{+}$. Definition 3.8 strengthens this property into a true martingale property. This corresponds to a generalization of the concept of fair pricing in the context of the benchmark approach, where a price process is said to be fair if it is a true martingale when denominated in units of the growth-optimal portfolio (see [51, Chapter 10] and [13, Section 2.2] for term structure models).
@@
Suppose that Assumption 3.1 holds and $\mathcal{D}\neq\emptyset$. For $D\in\mathcal{D}$ and $i,j\in I$ with $i<j$, assume that the term structures indexed by $i$ and $j$ are fairly priced by $D$, in the sense of Definition 3.8. Assume furthermore that Condition 3.7 holds and $S^{i}_{0}\leq S^{j}_{0}$. Then, it holds a.s. that
@@
The assumptions imply that the process $D(X^{0})^{-1}S^{k}B^{k}(\cdot,T)$ is a true martingale, for $k\in\{i,j\}$ and for all $T>0$. Therefore, making use of the terminal condition $B^{k}(T,T)=1$, it follows that
$S_{t}^{k}B^{k}(t,T)=\frac{\mathbb{E}[D_{T}(X_{T}^{0})^{-1}S_{T}^{k}|\mathcal{F}_{t}]}{D_{t}(X_{t}^{0})^{-1}}\quad\text{ a.s.,}$
@@ 3.4. The risk-neutral setup
We have so far developed a general framework which does not rely on the existence of a risk-neutral probability measure, being based on a weaker concept of market viability (see Section 2.2). Given the widespread use of risk-neutral valuation in finance, it is nevertheless useful to specialize Theorem 3.4 to the risk-neutral setup. Adopting the notation of Section 2.2, we say that a probability measure $\mathbb{Q}\sim\mathbb{P}$ on $(\Omega,\mathcal{F})$ is a risk-neutral probability if $X^{(i,T)}$ is a $\mathbb{Q}$-local martingale, for every $(i,T)\in I_{0}\times\mathbb{R}_{+}$. We recall from Girsanov’s theorem (see [37, Theorems III.3.11 and III.3.17]) that, if $\mathbb{Q}\sim\mathbb{P}$, there exist $\lambda\in L^{2}_{\rm loc}(W)$ and $\psi\in G_{\rm loc}(\mu)$ with $\psi>-1$ such that $W^{\mathbb{Q}}:=W+\int_{0}^{\cdot}\lambda_{s}{\rm d}s$ is a $d$-dimensional Brownian motion under $\mathbb{Q}$ and the compensator of $\mu$ under $\mathbb{Q}$ is given by
@@
Let $\mathbb{Q}$ be a probability measure on $(\Omega,\mathcal{F})$ with $\mathbb{Q}\sim\mathbb{P}$ and density process $D=(D_{t})_{t\geq 0}$, i.e., $D_{t}={\rm d}\mathbb{Q}|_{\mathcal{F}_{t}}/{\rm d}\mathbb{P}|_{\mathcal{F}_{t}}$, for all $t\geq 0$. [37, Lemma III.5.17] implies that $D$ admits a representation of the form (3.6). By definition, $\mathbb{Q}$ is a risk-neutral probability if and only if $D\in\mathcal{D}$. In view of Theorem 3.4, we only have to prove that (3.19)–(3.22) hold under $\mathbb{Q}$. For every $i\in I$, the process $Z^{i}$ can be represented as in (3.19) if and only if it is a special semimartingale under $\mathbb{Q}$ and this holds if and only if $c^{i}\mathbf{1}_{\{c^{i}>1\}}\ast\nu^{\mathbb{Q}}$ is of finite variation. Condition (3.5) for $t=T$ implies that $\int_{\{x\in E:c^{i}_{t}(x)>1\}}c^{i}_{t}(x)F^{\mathbb{Q}}_{t}({\rm d}x)<+\infty$ a.s. for a.e. $t\in\mathbb{R}_{+}$. Making use of condition (i) in Theorem 3.4 and the fact that $Z^{i}$ is a special semimartingale under $\mathbb{P}$, as a consequence of (3.1), we can compute
@@
    When applied to foreign exchange models (see Section 1.1.1), Corollary 3.12 extends the result of [46, Proposition 2.1.15] by requiring substantially weaker integrability properties.